Circular No. 6960

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1998 M6 J. Montani reports his discovery of a comet visually on the image-

display monitor during the course of routine scanning with the Spacewatch 0.9-m telescope at Kitt Peak. A hint of a very faint fan-shaped tail of length ∼ 10′′–20′′ lies in p.a. 175^{o}–200^{o}. The coma diameter measured 6′′ and 8′′ on June 30 and July 1, respectively. Available astrometry:

1998 UT  α_{2000}  δ_{2000}  m_{1}
June 30.34038  20^{h}57^{m}58.^{s}45  −17^{o}24′09′′.2  18.9
30.36233  20 57 57.44  −17 24 04.7  18.9
30.38417  20 57 56.42  −17 23 59.2  19.0
July 1.35790  20 57 11.23  −17 20 25.9  19.0
1.37000  20 57 10.63  −17 20 23.3  18.7
1.44301  20 57 07.19  −17 20 07.4  18.7
1.45423  20 57 06.64  −17 20 05.2  18.9

SUPERNOVA 1998cs IN UGC 10432 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Center for Astro-

physics, report: “A spectrum taken by K. Rines with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph) on July 1.2 UT shows SN 1998cs (cf. IAUC 6957) to be a type-Ia supernova before maximum light. Zabludoff et al. (1990, Ap.J. Suppl. 74) give a recession velocity for UGC 10432 of 9762 ± 43 km/s, which yields a high expansion velocity of 17 500 km/s (measured using the prominent Si II 635.5-nm feature).”

COMET C/1998 J1 (SOHO) Ephemeris extension to IAUC 6930 (orbital elements MPC 31893):

1998 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
July 6 8^{h}44.^{m}29 −41^{o}32.^{′}6 1.555 1.517 68.^{o}7 38.^{o}6  8.5
11 8 59.85 −43 27.5 1.654 1.610 69.5 36.3  8.8
16 9 14.88 −45 11.1 1.754 1.700 70.0 34.2  9.1
21 9 29.48 −46 45.7 1.854 1.788 70.2 32.3  9.4
26 9 43.70 −48 13.1 1.955 1.874 70.2 30.7  9.6
31 9 57.60 −49 34.8 2.056 1.958 70.1 29.2  9.8
Aug. 5 10 11.22 −50 51.8 2.157 2.041 69.7 27.8 10.0
10 10 24.60 −52 04.8 2.257 2.122 69.3 26.5 10.2
15 10 37.80 −53 14.5 2.358 2.202 68.6 25.4 10.4
1998 July 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green